The cost of 8 muffins and 2 quarts of milk is $18. The cost of 3 muffins and 1 quart of milk is $7.50. How much does 1 muffin and 1 quart of milk cost?
Answer by TimothyLamb(4379)

x = cost muffin
y = cost quart of milk
---
8x + 2y = 18
3x + y = 7.50
---
put the system of linear equations into standard form
---
8x + 2y = 18
3x + y = 7.50
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = cost muffin = 1.50
y = cost quart of milk = 3.00
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php